### v3

@@ 4.3.2 Discrepancy-Based Alignment
Discrepancy-based deep domain adaptation (DA) methods aim to reduce the divergence between marginal and/or conditional distributions of source and target domains by incorporating distribution alignment metrics—such as Maximum Mean Discrepancy (MMD)—into deep neural network architectures [108]. These adaptation strategies typically employ additional alignment layers to encourage the learning of domain-invariant, task-relevant representations, as illustrated in Figure 13.
@@
Long et al. [109] pioneered the Deep Adaptation Network (DAN), which extends convolutional neural networks to domain adaptation by aligning the distributions of learned features across domains. Baktashmotlagh et al. [110] proposed a domain-invariant projection approach using MMD, while Shen et al. [111] introduced a Wasserstein Distance Guided Representation Learning (WDGRL) strategy. Sun et al. [112] extended the CORAL framework by proposing Deep CORAL, which aligns activation correlations in deep layers through nonlinear transformations.
In the context of agriculture, Ferreira [113] proposed an Unsupervised Domain Adaptation framework based on vision Transformers (UDA-SegFormer) to improve crop row and gap detection across farms. By aligning domain-specific features and simplifying pseudo-labels, the model achieved superior generalization performance in challenging environments (e.g., shadows, curved rows), outperforming traditional CNN-based networks such as PSPNet and DeepLabV3+. This work represents the first integration of Transformers with UDA in agricultural settings, enabling label-efficient and robust crop monitoring.
Takahashi et al. [34] enhanced agricultural image recognition by performing category-wise feature space alignment. Building on DAN, Long et al. [114] proposed the Multi-Representation Adaptation Network (MRAN) [115] and the Deep Sub-domain Adaptation Network (DSAN) [116], which leverage sub-domain partitioning and multiple feature representations to address inter-domain variability.
Real-world high-resolution satellite (HSR) images often differ across regions in terms of scale, resolution, and spectral style [117]. To address such heterogeneity, Zhu et al. [118] proposed the Weak Pseudo-supervised Decorrelation Sub-domain Adaptation (WPS-DSA) network for land use classification, which effectively handles spatial and temporal feature distribution shifts.
In the field of cotton boll status recognition, Li [119] introduced NCADA, an unsupervised DA method based on Neighborhood Component Analysis, along with the first benchmark dataset for in-field cotton boll imagery. NCADA outperformed traditional classification methods and achieved competitive accuracy under real-world distribution shifts, paving the way for automated phenotyping in precision agriculture.
@@
For hyperspectral image (HSI) domain adaptation, Garea et al. [120] introduced TCANet, which incorporates Transfer Component Analysis (TCA) into a deep architecture. Ian et al. [77] developed DDA-Net to enable unsupervised knowledge transfer across HSI datasets. Li et al. [121] proposed a two-stage deep domain adaptation (TDDA) framework: Stage I minimizes MMD-based domain shifts via deep embedding; Stage II employs a spatial-spectral twin network to extract discriminative features. Zhang et al. [122] presented TST-Net, which leverages graph convolutional networks (GCNs) and convolutional layers to encode topological and semantic information for cross-scene classification. Wang et al. [123] introduced a GNN-based DA model that jointly aligns domain and class distributions through CORAL integration. Liang et al. [124] proposed AMF-FSL, an attention-driven few-shot learning framework for HSI classification, which employs target-class alignment, domain attention, and multi-source fusion modules.
@@
In plant pathology, Yan et al. [125] designed a hybrid sub-domain alignment strategy for cross-species plant disease severity recognition. Fuentes et al. [126] developed a domain-adaptive classification model for tomato diseases that accommodates environmental variability. However, these approaches fall short in addressing the complexities of cross-species disease diagnosis. To this end, Wu [127] introduced MSUN, a multi-representation sub-domain alignment network with uncertainty regularization. MSUN effectively handles intra-domain ambiguity, inter-class overlap, and inter-domain heterogeneity, which are critical challenges in plant disease recognition under cross-species settings.
@@
Othman et al. [128] proposed a domain adaptation (DA) network for cross-scene classification, utilizing a two-stage strategy of pretraining and fine-tuning. Their method ensures accurate classification of source domain samples while aligning feature distributions between domains and preserving the geometric structure of the target domain data. Similarly, Lu et al. [129] introduced a Multi-Source Compensation Network (MSCN) designed for cross-scene classification. This architecture incorporates both cross-domain alignment modules and classifier compensation modules, which jointly reduce domain shift and align class semantics across multiple source domains. Zhu et al. [130] developed the Attention-based Multi-Scale Residual Adaptation Network (AMRAN), which combines a residual adaptation module for marginal distribution alignment, an attention module for enhanced feature extraction, and a multi-scale adaptation module that targets conditional distribution alignment. In the SAR image domain, Geng et al. [131] proposed the Deep Joint Distribution Adaptation Network (DJDAN), which simultaneously addresses both marginal and conditional distribution discrepancies for effective transfer learning in SAR image classification tasks.
@@
In recent years, self-supervised learning (SSL) has emerged as a powerful paradigm in the absence of labeled data, demonstrating strong capability in feature representation learning [132]. Zhao et al. [133] proposed a Contrastive Learning and Domain Adaptation Layer (CLA) framework for leaf disease recognition, which innovatively combines self-supervised pretraining with fine-tuned domain adaptation. This two-stage approach is particularly effective in agricultural contexts with large-scale unlabeled image collections.
@@
In the self-supervised pretraining stage, CLA employs contrastive learning to train an encoder on a large pool of unlabeled and distribution-confounded images. The goal is to maximize similarity between augmented views of the same image (positive pairs) while minimizing similarity across different images (negative pairs), akin to the contrastive loss used in SimCLR.
@@
In the second stage, CLA fine-tunes the learned representations with a limited amount of labeled data using a Domain Adaptation Layer (DAL), which jointly optimizes supervised loss and domain alignment loss. The loss function is defined as:
$L_{DAL}=L_{\text{supervised}}+\lambda\cdot L_{\text{MMD}}$ (7)
@@
where $L_{\text{supervised}}$ denotes the conventional classification loss on labeled data, $L_{\text{MMD}}$ represents the domain alignment loss based on Maximum Mean Discrepancy (MMD), and $\lambda$ is a hyperparameter that balances the two objectives.
@@
Experimental results demonstrated that CLA significantly outperformed baseline and comparison methods in both domain alignment and final classification accuracy, achieving a peak accuracy of 90.52%. Further ablation studies and sensitivity analyses were conducted to identify key contributing factors, providing deeper insights into the underlying mechanisms.
@@
The CLA framework provides a compelling example of combining self-supervised contrastive learning with domain adaptation to improve generalization performance in leaf disease recognition under low-label regimes. Its capability to leverage unlabeled data while mitigating domain shifts holds strong potential for broad adoption in agricultural image analysis. By enhancing robustness to domain variance and environmental noise, CLA offers a scalable and label-efficient solution for intelligent plant health monitoring systems.
@@ 4.4 Summary of Deep Domain Adaptation (DA) Methods
With the rapid development of deep learning technologies, deep domain adaptation (DA) methods have demonstrated significant advantages in agricultural image analysis. These methods, utilizing deep neural networks (DNNs) and adversarial training, are capable of automatically learning domain-invariant feature representations, effectively addressing the complex problem of domain shift. Typically, deep DA methods are classified into supervised, semi-supervised, and unsupervised categories.
In supervised deep DA methods, source domain classification loss and cross-domain alignment loss are jointly optimized to align the feature distributions between the source and target domains, thereby improving the model’s generalization ability in the target domain. Semi-supervised deep DA methods combine fully labeled data from the source domain with a small amount of labeled data from the target domain, balancing labeling costs with model performance, making them particularly suitable for agricultural scenarios with limited labeled data. Unsupervised deep DA methods rely entirely on labeled data from the source domain and unlabeled data from the target domain, achieving domain adaptation through feature alignment or generative models. These methods have a wide range of applications.
The primary advantage of deep DA methods lies in their ability to handle high-dimensional, complex agricultural data and, through end-to-end learning, automatically extract deep features. This enables them to effectively mitigate domain shifts caused by differences in sensors, environmental changes, and crop growth stages. However, deep DA methods also present certain challenges, including high computational resource demands, long model training times, and poor model interpretability. These issues can be limiting factors in agricultural environments with limited resources or high real-time performance requirements. Nonetheless, with continuous technological advancements, deep DA methods are poised to play an increasingly important role in precision agriculture and smart farming.
@@ 5 Experimental Results: Evaluation, Comparison, and Analysis
5.1 Evaluation Metrics
We adopt a diverse set of evaluation metrics to comprehensively assess the performance of classification, regression, and segmentation models in agricultural image analysis. The definitions and corresponding formulas are listed below.
Average Accuracy (AA)
$\text{AA}=\frac{1}{C}\sum_{i=1}^{C}\frac{N_{ii}}{\sum_{j=1}^{C}N_{ij}}$ (8)
Where: $C$ is the number of classes, and $N_{ij}$ denotes the number of samples whose true label is class $i$ and predicted as class $j$. This metric gives equal weight to all classes regardless of class imbalance.
Root Mean Squared Error (RMSE)
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (9)
Where: $y_{i}$ is the ground truth value, and $\hat{y}_{i}$ is the predicted value. RMSE expresses prediction error in the same units as the target variable and is widely used in regression tasks.
Coefficient of Determination ($R^{2}$)
Where: $\bar{y}$ denotes the mean of the ground truth values. The $R^{2}$ score measures the proportion of variance in the dependent variable that is predictable from the independent variables.
Intersection over Union (IoU)
$\text{IoU}=\frac{|P\cap G|}{|P\cup G|}$ (11)
Where: $P$ is the predicted region and $G$ is the ground truth region. IoU is a widely used metric for evaluating semantic segmentation performance.
Completeness (COM)
$\text{COM}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (12)
COM quantifies the proportion of actual agricultural land pixels that were correctly identified.
Correctness (COR)
$\text{COR}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (13)
COR reflects the proportion of predicted agricultural land pixels that are truly correct.
Segmentation F1 Score
[…]
This F1 score variant harmonizes completeness and correctness to evaluate the quality of segmentation results.
@@ 5.1 Evaluation Metrics
NASS N/A Tabular data Multi-year regionwise N/A N/A
@@ 5.2 Plant Disease Detection Experiments
To evaluate domain adaptation for plant disease detection, we conducted four sets of experiments using five benchmark datasets: PlantVillage (PVD) [134], PlantDoc [135], Plant-Pathology [136], Corn-Leaf-Diseases, and Tomato-Leaf-Diseases. Among them, PlantVillage (PVD) was consistently used as the source domain, as it features a controlled acquisition environment with clean backgrounds that facilitates model training.
@@
Table 3 summarizes the results of eight models—Baseline, DAAN, D-CORAL, DAN, DANN, MRAN, DSAN, and MSUN—on the four experiments: C-PD, PVD-PP, PVD-CLD, and PVD-TLD. Compared with the multi-species transfer scenario in C-PD, single-species domain adaptation (e.g., PVD-CLD and PVD-TLD) generally performs better due to clearer intra-class boundaries. MSUN consistently achieves top performance across most tasks, benefiting from its multi-representation modeling and subdomain alignment strategies that address high inter-class similarity and intra-class variability.
@@
In the plant disease transfer detection task, MSUN demonstrates significant advantages across a range of experimental configurations. Whether under multi-plant mixed transfer (C-PD) or single-plant transfer (PVD-PP, PVD-CLD, PVD-TLD), MSUN consistently achieves superior average accuracy compared to mainstream domain adaptation methods.
Specifically, in the C-PD task, MSUN achieves an average accuracy of 56.06%, outperforming the baseline model by more than 25 percentage points. It also surpasses advanced methods such as DSAN and MRAN, illustrating its strong adaptability to large domain shifts and complex inter-class relationships.
In the PVD-PP experiment, MSUN obtains the highest accuracy across all three disease categories, reaching an average accuracy of 72.31%. It significantly outperforms the second-best method MRAN (67.45%) and the traditional DANN (50.04%), reflecting its enhanced transferability in single-plant scenarios.
In the PVD-CLD setup, MSUN achieves an impressive average accuracy of 96.78%, particularly excelling in the “blight” (93.81%) and “healthy” (84.77%) categories. This highlights MSUN’s strong ability to capture discriminative features and reduce inter-class confusion.
The PVD-TLD task, featuring fine-grained categories and highly imbalanced distributions, presents greater challenges. While most methods perform poorly, MSUN still attains an average accuracy of 50.58%. Notably, it achieves 80.79% in the low-sample “Septoria spot” category, demonstrating superior generalization to minority classes and robustness to data imbalance.
Overall, MSUN exhibits consistent and leading performance across various plant disease transfer tasks, validating the effectiveness of its multi-representation learning and subdomain alignment strategy. Compared to traditional domain adaptation approaches, MSUN not only improves classification accuracy but also offers enhanced robustness and generalization in the face of class imbalance, multi-disease complexity, and cross-plant variability.

### v4

@@ 4.3.2 Discrepancy-Based Alignment
Discrepancy-based deep domain adaptation (DA) methods aim to reduce the divergence between marginal and/or conditional distributions of the source and target domains by incorporating distribution alignment metrics—such as Maximum Mean Discrepancy (MMD)—into deep neural network architectures [109]. These strategies typically employ additional alignment layers to encourage the learning of domain-invariant, task-relevant representations, as illustrated in Figure 13.
@@
Long et al. [110] pioneered the Deep Adaptation Network (DAN), which extends convolutional neural networks for domain adaptation by aligning the distributions of learned features across domains. Baktashmotlagh et al. [111] proposed a domain-invariant projection method using MMD, while Shen et al. [112] introduced the Wasserstein Distance Guided Representation Learning (WDGRL) framework. Sun et al. [113] extended the CORAL framework with Deep CORAL, aligning activation correlations in deep layers through nonlinear transformations.
In agriculture, Ferreira [114] proposed an unsupervised domain adaptation framework based on vision Transformers (UDA-SegFormer) to enhance crop row and gap detection across farms. By aligning domain-specific features and simplifying pseudo-labels, the model achieved superior generalization in challenging environments (e.g., shadows, curved rows), outperforming traditional CNN-based networks such as PSPNet and DeepLabV3+. This work represents the first integration of Transformers with UDA in agricultural contexts, enabling label-efficient and robust crop monitoring.
Guo et al. [115] propose masked image consistency and multi-granularity alignment to enhance domain adaptation for plant disease detection. Their method reduces domain shift between source and target data, improving cross-domain robustness. This approach achieves state-of-the-art performance in agricultural disease recognition.Gao et al. [116] develop a cross-domain transfer learning framework for weed segmentation using both ground and UAV images. Their model bridges domain gaps between aerial and ground-level perspectives in precision farming. The solution enables accurate weed mapping for targeted agricultural management.Fu et al. [117] introduce a multi-domain data augmentation strategy combined with isolated test-time adaptation for crop disease detection. This dual approach enhances model generalization to unseen field conditions without source data access. Their method significantly improves robustness against domain shifts.
Takahashi et al. [35] enhanced agricultural image recognition via category-wise feature space alignment. Building on DAN, Long et al. [118] proposed the Multi-Representation Adaptation Network (MRAN) [119] and the Deep Sub-domain Adaptation Network (DSAN) [120], which utilize sub-domain partitioning and multi-representation learning to better address inter-domain variability.
Real-world high-resolution satellite (HRS) imagery often varies across regions in terms of scale, resolution, and spectral characteristics [121]. To address this heterogeneity, Zhu et al. [122] proposed the Weak Pseudo-supervised Decorrelation Sub-domain Adaptation (WPS-DSA) network for land use classification, which effectively handles both spatial and temporal feature distribution shifts.
In the field of cotton boll status recognition, Li [123] introduced NCADA, an unsupervised DA method based on Neighborhood Component Analysis, along with the first benchmark dataset for in-field cotton boll imagery, as shown in Figure 14. NCADA outperformed traditional classification methods and achieved competitive accuracy under real-world distribution shifts, paving the way for automated phenotyping in precision agriculture.
@@
For hyperspectral image (HSI) domain adaptation, Garea et al. [124] introduced TCANet, which integrates Transfer Component Analysis (TCA) into a deep network. Ian et al. [78] developed DDA-Net to facilitate unsupervised knowledge transfer across HSI datasets. Li et al. [125] proposed a two-stage deep domain adaptation (TDDA) framework: Stage I minimizes MMD-based domain shifts via deep embedding; Stage II employs a spatial-spectral twin network to extract discriminative features. Zhang et al. [126] presented TST-Net, which combines graph convolutional networks (GCNs) and convolutional layers to encode both topological and semantic information for cross-scene classification. Wang et al. [127] proposed a GNN-based DA model that jointly aligns domain and class distributions through CORAL integration. Liang et al. [128] proposed AMF-FSL, an attention-driven few-shot learning framework for HSI classification that incorporates target-class alignment, domain attention, and multi-source fusion modules.
@@
In plant pathology, Yan et al. [129] designed a hybrid sub-domain alignment strategy for cross-species plant disease severity recognition. Fuentes et al. [130] developed a domain-adaptive classification model for tomato diseases that accounts for environmental variability. However, these approaches often fail to address the full complexity of cross-species disease diagnosis. To this end, Wu [131] proposed MSUN, a multi-representation sub-domain alignment network with uncertainty regularization. MSUN effectively addresses intra-domain ambiguity, inter-class overlap, and inter-domain heterogeneity—key challenges in plant disease recognition across species.
@@
Othman et al. [132] introduced a DA network for cross-scene classification, utilizing a two-stage training strategy combining pretraining and fine-tuning. Their approach ensures accurate classification of source-domain samples while aligning feature distributions between domains and preserving the geometric structure of target-domain data. Similarly, Lu et al. [133] proposed the Multi-Source Compensation Network (MSCN), which incorporates both cross-domain alignment modules and classifier compensation mechanisms to jointly reduce domain shift and align class semantics across multiple source domains. Zhu et al. [134] developed the Attention-based Multi-Scale Residual Adaptation Network (AMRAN), which includes a residual adaptation module for marginal distribution alignment, an attention module for feature enhancement, and a multi-scale adaptation module for conditional distribution alignment. In the SAR image domain, Geng et al. [135] proposed the Deep Joint Distribution Adaptation Network (DJDAN), which simultaneously aligns marginal and conditional distributions to facilitate transfer learning in SAR image classification.
@@
Recently, self-supervised learning (SSL) has emerged as a powerful paradigm in the absence of labeled data, offering strong capabilities in feature representation learning [136]. Zhao et al. [137] proposed the Contrastive Learning and Domain Adaptation Layer (CLA) framework for leaf disease recognition, which innovatively combines self-supervised pretraining with fine-tuned domain adaptation. This two-stage approach is particularly effective for agricultural scenarios with large-scale unlabeled datasets.
@@
In the self-supervised pretraining stage, CLA employs contrastive learning to train an encoder on a large pool of unlabeled, domain-confounded images. The objective is to maximize the similarity between augmented views of the same image (positive pairs) while minimizing similarity between different images (negative pairs), similar to the contrastive loss used in SimCLR.
@@
In the second stage, CLA fine-tunes the learned representations with limited labeled data using a Domain Adaptation Layer (DAL), which jointly optimizes supervised loss and domain alignment loss. The total loss function is defined as:
$L_{\text{DAL}}=L_{\text{supervised}}+\lambda\cdot L_{\text{MMD}}$ (7)
@@
Where: $L_{\text{supervised}}$ is the classification loss on labeled data, $L_{\text{MMD}}$ is the Maximum Mean Discrepancy-based alignment loss, and $\lambda$ is a trade-off hyperparameter.
@@
Experimental results show that CLA significantly outperforms baseline and comparison models in both domain alignment and final classification accuracy, achieving a peak accuracy of 90.52%. Further ablation and sensitivity analyses reveal key contributing factors, offering deeper insights into its mechanism.
@@
The CLA framework exemplifies the effective combination of self-supervised contrastive learning and domain adaptation to enhance generalization performance in leaf disease recognition under low-label regimes. Its ability to leverage unlabeled data while mitigating domain shift makes it a scalable and label-efficient solution for intelligent plant health monitoring systems.
@@ 4.4 Summary of Deep Domain Adaptation (DA) Methods
With the rapid advancement of deep learning technologies, deep domain adaptation (DA) methods have demonstrated significant advantages in agricultural image analysis. These approaches, based on deep neural networks (DNNs) and adversarial training, are capable of automatically learning domain-invariant feature representations, effectively addressing the challenges posed by domain shift. Deep DA methods are typically categorized into supervised, semi-supervised, and unsupervised frameworks.
In supervised deep DA methods, the source-domain classification loss and cross-domain alignment loss are jointly optimized to align the feature distributions between source and target domains. This enhances the model’s generalization capability in the target domain. Semi-supervised deep DA methods utilize fully labeled source-domain data along with a small portion of labeled target-domain data. By balancing annotation costs and model performance, these methods are particularly suitable for agricultural scenarios where labeled target data is scarce.
Unsupervised deep DA methods rely exclusively on labeled data from the source domain and unlabeled data from the target domain. Domain adaptation is achieved through feature alignment strategies (e.g., adversarial learning, distribution matching) or generative models (e.g., GAN-based feature synthesis), making them widely applicable in real-world agricultural tasks.
The key advantage of deep DA methods lies in their ability to process high-dimensional, complex agricultural data and extract domain-invariant representations via end-to-end training. This enables effective mitigation of domain discrepancies caused by sensor variability, environmental heterogeneity, and crop phenological stages. However, these methods also face several challenges, such as high computational demands, prolonged training times, and limited interpretability. These drawbacks may constrain their deployment in resource-limited environments or scenarios requiring real-time performance.
Despite these limitations, ongoing advancements in model efficiency, interpretability, and hardware acceleration continue to expand the potential of deep DA methods. They are expected to play an increasingly critical role in the development of precision agriculture and intelligent farming systems.
@@ 5 Experimental Results: Evaluation, Comparison, and Analysis
5.1 Datasets Overview
The experiments conducted in this study utilize several prominent agricultural image datasets to evaluate domain adaptation techniques effectively. Specifically, the remote sensing datasets include:
  • DeepGlobe: Primarily employed for urban and rural land-cover classification and segmentation. Download via Dataset Ninja or official DeepGlobe 2018 Challenge website.
  • GID: Used extensively for high-resolution land-use and land-cover mapping. Official site: https://captain-whu.github.io/GID/, with OneDrive/Baidu download links.
  • LoveDA: Designed for detailed semantic segmentation tasks including agricultural and geographic information extraction. GitHub repository at https://github.com/Junjue-Wang/LoveDA, with dataset hosted on Zenodo.
For crop disease detection and classification tasks, the datasets employed are:
  • PlantVillage: Widely used for benchmarking plant disease classification methods under controlled laboratory conditions. Available on Kaggle: https://www.kaggle.com/datasets/emmarex/plantdisease and GitHub.
  • PlantDoc: Targeted toward real-world plant disease diagnosis under diverse environmental conditions. GitHub repository (PlantDoc-Dataset) and listed on Papers With Code.
  • Plant-Pathology: Employed for identification and severity assessment of foliar diseases. Provided through the Kaggle FGVC8 competition.
  • Corn-Leaf-Diseases: Specifically developed for maize disease detection. Available on Kaggle as “Corn or Maize Leaf Disease Dataset”.
  • Tomato-Leaf-Diseases: Focused on tomato leaf disease monitoring. Downloadable from Kaggle or Mendeley Data.
Additionally, the following dataset provides tabular agricultural data:
  • NASS: Widely applied in crop yield prediction, agricultural policy analysis, and cross-domain productivity modeling. Available from the official USDA NASS portal: https://www.nass.usda.gov.
A summary of these datasets—including spatial resolution, sensor type, image dimensions, data partitioning, and their respective application scenarios—is presented in Table 2. Collectively, these datasets represent a comprehensive range of agricultural scenarios, sensor modalities, and domain challenges, thereby serving as robust benchmarks for evaluating domain adaptation approaches in agricultural image analysis.
@@ 5.1 Datasets Overview
NASS N/A Tabular Data Multi-year (regionwise) N/A N/A
5.2 Evaluation Metrics
We adopt a diverse set of evaluation metrics to comprehensively assess the performance of classification, regression, and segmentation models in agricultural image analysis. The definitions and corresponding formulas are provided below.
Average Accuracy (AA)
$\text{AA}=\frac{1}{C}\sum_{i=1}^{C}\frac{N_{ii}}{\sum_{j=1}^{C}N_{ij}}$ (8)
Where: $C$ is the number of classes, and $N_{ij}$ denotes the number of samples with true label $i$ predicted as class $j$. This metric assigns equal weight to each class, mitigating the effects of class imbalance.
Root Mean Squared Error (RMSE)
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (9)
Where: $y_{i}$ is the ground truth value, and $\hat{y}_{i}$ is the predicted value. RMSE expresses the prediction error in the same units as the target variable and is widely used in regression tasks.
Coefficient of Determination ($R^{2}$)
Where: $\bar{y}$ is the mean of the ground truth values. The $R^{2}$ score quantifies the proportion of variance in the dependent variable that is predictable from the independent variables.
Intersection over Union (IoU)
$\text{IoU}=\frac{|P\cap G|}{|P\cup G|}$ (11)
Where: $P$ is the predicted region and $G$ is the ground truth region. IoU is a standard metric for evaluating semantic segmentation performance.
Completeness (COM)
$\text{COM}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (12)
COM measures the proportion of actual agricultural land pixels that were correctly identified by the model.
Correctness (COR)
$\text{COR}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (13)
COR indicates the proportion of predicted agricultural land pixels that are truly correct.
Segmentation F1 Score
$\text{F1}_{\text{seg}}=\frac{2\cdot\text{COM}\cdot\text{COR}}{\text{COM}+\text{COR}}$ (14)
This segmentation-specific F1 score harmonizes completeness and correctness to evaluate overall segmentation quality.
@@ 5.3 Plant Disease Detection Experiments
To evaluate domain adaptation for plant disease detection, we conducted four sets of experiments using five benchmark datasets: PlantVillage (PVD) [138], PlantDoc [139], Plant-Pathology [140], Corn-Leaf-Diseases, and Tomato-Leaf-Diseases, as shown in Figure 15. Among them, PlantVillage (PVD) was consistently used as the source domain due to its controlled acquisition environment and clean backgrounds, which facilitate robust model training.
@@
Table 3 summarizes the results of eight models—Baseline, DAAN, D-CORAL, DAN, DANN, MRAN, DSAN, and MSUN—on four transfer tasks: C-PD, PVD-PP, PVD-CLD, and PVD-TLD. Compared to the multi-crop mixed transfer scenario in C-PD, single-crop domain adaptation tasks (e.g., PVD-CLD and PVD-TLD) generally yield better performance due to more consistent intra-class boundaries. MSUN consistently achieves top accuracy across all settings, owing to its multi-representation learning and subdomain alignment mechanisms, which address high inter-class similarity and intra-class variability.
@@
In the plant disease transfer detection task, MSUN demonstrates significant advantages across diverse experimental configurations. Whether in multi-crop mixed transfer (C-PD) or single-crop transfer (PVD-PP, PVD-CLD, PVD-TLD), MSUN consistently achieves superior average accuracy compared to state-of-the-art domain adaptation methods.
Specifically, in the C-PD task, MSUN achieves an average accuracy of 56.06%, outperforming the baseline by over 25 percentage points. It also surpasses advanced models such as DSAN and MRAN, showcasing strong adaptability to substantial domain shifts and complex inter-class relationships.
In the PVD-PP experiment, MSUN attains the highest accuracy across all three disease categories, reaching an average of 72.31%. It significantly outperforms MRAN (67.45%) and DANN (50.04%), highlighting its enhanced transferability in single-crop scenarios.
Under the PVD-CLD setting, MSUN achieves an outstanding average accuracy of 96.78%, particularly excelling in the “Blight” (93.81%) and “Healthy” (84.77%) classes. This underscores its strong capability in extracting discriminative features and reducing inter-class confusion.
The PVD-TLD task, characterized by fine-grained classes and severe data imbalance, poses greater challenges. While most methods show limited effectiveness, MSUN still achieves an average accuracy of 50.58%. Notably, it attains 80.79% accuracy in the low-sample “Septoria spot” category, indicating superior generalization to minority classes and robustness against sample imbalance.
Overall, MSUN exhibits consistent and leading performance across plant disease transfer tasks, validating the effectiveness of its multi-representation learning and subdomain alignment strategies. Compared to traditional domain adaptation approaches, MSUN not only enhances classification accuracy but also improves robustness and generalization under conditions of class imbalance, multi-disease complexity, and cross-crop variability.